a freight plane has a total mass o 4.00 x 10^5 kg (including cargo). It starts from rest at sea level and within 4.00 minutes gets to a speed of 90.0 m/s at a height of 2.00 km. Ignorining the air drag force and friction, what is the averaged required power of its engines?
Answer #1

total energy gained

E = mgh + 0.5 m v^2

E = 4* 10^5* ( 9.8 * 2000 + 0.5* 90^2)

E = 9.46* 10^9 N

power = energy gained / time

P = 9.46* 10^9 / ( 4* 60)

P = 3.942* 10^7

=======

Comment before rate in case any doubt, will reply for sure.. goodluck